# Local wind speed forecasting at short time horizons relying on both Numerical Weather Prediction and observations from surrounding stations

### v1

@@ 3.2 A neural network model designed to handle station data and NWP predictions
Our “hybrid” neural network model that integrates NWP short-term forecasts with local ground station data is designed to process 3 distinct sets of MeteoNet input data: stations data, AROME predictions, and ARPEGE predictions. It is illustrated in Fig. 2. One can see that the model is composed of 3 branches devoted to handle the different input features. These features, associated with each branch, are summarized in Table 1. Let us describe more precisely the input feature data associated with each of the 3 parts of the model for any given location, that is, for any site referenced as a “ground station” in MeteoNet dataset.
@@
Table 1: Summary of the MeteoNet input data used in the model ${\mathcal{M}}$
@@
Let ${\bf{GS}}_{d,t}^{s}$,${\bf{AR}}_{d,t}^{s}$ and ${\bf{AP}}_{d,t}^{s}$ be the feature tensors built from respectively the ground station, AROME and ARPEGE data at some given day $d$, time $t$ and corresponding to some given site $s$. The first one, associated with weather ground station data, is composed of 3 different variables: the wind speed data ($u$,$v$ components) and surface temperature data. Other measures like pressure or humidity have drastically lower number of available data, so we chose to restrict ourselves to wind speed and temperature features. For each site, we also include these 3 variables measured at the 10 closest surrounding weather stations for which the data are also available. In that respect, at any given $(d,t,s)$, the number of station input variables is $N_{C}=33$. At a given current time $t$, we are considering not only the measures at $t$ but also the measures obtained 6 hours in the past i.e. at times $t-1,\ldots,t-6$. At this stage, it is worth noting that the MeteoNet dataset provides ground station observations at a sampling frequency of 6 minutes. However, our experiments have shown that increasing the time resolution does not lead to any significant improvement in forecasting performance within our model. Therefore, to maintain model simplicity and minimize the number of parameters, we restrict our analysis to data at an hourly scale. The number considered of time periods (past or present) is then $N_{H}=7$, so that the overall ${\bf{GS}}_{d,t}^{s}$ tensor shape is $(7\times 33)$.
@@
Among all AROME prediction data, we select wind eastward and northward components at 10 $m$, mean-sea-level pressure, 2 m temperature and relative humidity so that the number of input variables is $N_{C}=5$. We keep only data around the considered ground station position, i.e., we consider the set of $11\times 1$ 11 AROME grid points centered around the ground station, namely if $i_{0},j_{0}$ are the coordinates of the grid point closest to the target ground station $s$, then the considered AROME predictions are those within a spatial extent of ±$\pm 0.125^{\circ}$ latitude/longitude around this grid point. Moreover, we keep all prediction horizons $t+1,t+2,\ldots,t+6$ so that $N_{H}=6$. The shape of the tensors ${\bf{AR}}^{s}_{d,t}$ is therefore $6\times 5\times 11\times 11$. We note that, since AROME forecasts data are only available for $0\leq t\leq 24$, available features are restricted by the choice of the largest horizon $h_{H}$. That is, for each day $d$, we take $t$ to be the range $t=0,1,\ldots,18$.
@@
The ${\bf{AP}}^{s}_{d,t}$ tensor, which contains the data from the ARPEGE NWP model, is constructed along the same line: we keep only wind eastward and wind northward components together with temperature at all 7 isobar levels and we also consider the pressure levels at all 7 heights. As far as prediction horizons are concerned, we consider the same 6 next hours set of horizons as for AROME. When the horizon is not available, we consider the closest horizon multiple of 3 hours (indeed, ARPEGE predictions are not available on a regular time grid, as the time step varies from 1 hour in the morning to 3 hours in the afternoon). Concerning the grid size, we use all grid points within an interval ±$\pm 0.2^{\circ}$ around the closest point to the station. Notice that this spatial extent is larger than former AROME one but, due to the lower resolution of ARPEGE, it corresponds to a smaller ($5\times 5$) grid size. Accordingly, the ARPEGE input tensors have a shape $6\times 4\times 7\times 5\times 5$.
@@
As represented in Fig. 2 by blue 3D boxes, the branches that process AROME data (${\bf{AR}}^{s}_{d,t}$) and ARPEGE data (${\bf{AP}}^{s}_{d,t}$) inputs begin with a time-distributed convolution which applies the same spatial (2D for AROME and 3D for ARPEGE) convolution operation independently to each time step in a sequence of inputs, treating each time step as a separate sample. This allows the model to extract spatial features along the temporal axis while maintaining the temporal dimension for subsequent processing. For each of the $6$ times indices, the output of convolution layer is then flattened in order to provide a 2 dimensional tensor (red circular boxes on the sketch, of shape $(6,816)$ and (6,208)6,208)6 , 208 ) for AROME and ARPEGE branches respectively), such that they have the same number of dimension as the Station branch (whose input shape is $(7,33)$). In each of the 3 branches, a set of temporal and static features are also provided at each time step $t$ by tensor ${\bf{C}}^{s}_{d,t}$. These values are the time of the day $t$, the day of the year $d$ and the location (latitude, longitude and altitude) of the station $s$. ${\bf{C}}^{s}_{d,t}$ is then fed to a fully connected layer which encodes this information within a 16-dimensional tensor. For the station branch, a similar encoding of the position of each of the 10 neighboring sites, ${\bf{D}}^{s}$ with respect to $s$ is also provided. These tensors are then repeated at each of the input times (7 times for station data and 6 times for AROME and ARPEGE branches) and then concatenated to former input tensors or outputs of convolution layers.
As illustrated by green rectangles in Fig. 2, the so-obtained tensors are then processed through 2 or 3 stacked “Long Short-Term Memory” (LSTM) layers, a recurrent neural network architecture that is widely used to process sequential information (Goodfellow et al., 2016). The last outputs are finally entered into a fully connected layer. The outputs of each branch are finally concatenated into a single tensor to be processed by a final fully connected layer and then by the output layer, which is also a fully connected layer with a rectified linear unit (ReLU) activation that returns a vector of 6 positive values representing the predicted wind speeds at horizons $h=1,2,\ldots,6$ hours. Hereafter we refer to such a model with 3 distinct branches separately devoted to leverage ground station, AROME and ARPEGE data, as the ${\mathcal{M}}$ model. Within the notations of Eqs. (2) and (3), one has,
${\widehat{{\bf{V}}}}^{s}_{d,t+h}={\mathcal{M}}({\bf{X}}^{s}_{d,t},{\bf{\Theta}})\;$ (5)
@@
with ${\bf{X}}^{s}_{d,t}=\Big{[}{\bf{GS}}^{s}_{d,t},{\bf{AR}}^{s}_{d,t},{\bf{AP}}^{s}_{d,t},{\bf{C}}^{s}_{d,t},{\bf{D}}^{s}\Big{]}$. We will also consider the restriction of the model to a single type of input, i.e., one that considers individually ${\bf{GS}}^{s}_{t}$, ${\bf{AR}}^{s}_{t}$ or ${\bf{AP}}^{s}_{t}$ and their corresponding branch. When ${\mathcal{M}}$ is reduced to its “ground station” “AROME” or its “ARPEGE” branch, the corresponding models will be denoted as respectively ${\mathcal{M}}_{GS}$ model, ${\mathcal{M}}_{AR}$ model and ${\mathcal{M}}_{AP}$ model. Regarding the number of parameters to be learned in each model, ${\mathcal{M}}$, as previously described, consists of 262k parameters, while ${\mathcal{M}}_{AR}$, ${\mathcal{M}}_{AP}$, and ${\mathcal{M}}_{GS}$ contain 138k, 84k, and 39k parameters, respectively.
Let us finally notice that, as remarked previously, one can build ${\mathcal{M}}^{\prime}$, a variant of ${\mathcal{M}}$, that outputs the wind speed probability distribution instead of its expected value:
${\widehat{\rho}}^{s}_{d,t+h}={\mathcal{M}}^{\prime}({\bf{X}}^{s}_{d,t},{\bf{\Theta}})\;$ (6)
In this case, the last layer of ${\mathcal{M}}^{\prime}$ is a fully connected layer which outputs a $(n,6)$ tensor, where $n$ is the number of parameters necessary to define the chosen probability distribution ${\widehat{\rho}}^{s}_{d,t+h}$. In this work, we choose a law with $n=3$ parameters the so-called “Multifractal-Rice” (M-Rice) distribution, which has proven to be particularly suitable for wind speed forecasting when compared to other distributions (Baggio and Muzy, 2024). The M-Rice distribution has been introduced in Baïle et al. (2011) where the authors considered, inspired by the random cascade picture of fully developed turbulence, a Rice-distribution with a random scale parameter that follows, e.g., a log-normal distribution. Its density involves 3 parameters $(\nu,\sigma^{2},\lambda^{2}$). More details on this distribution and on the numerical method used to evaluate it are given in the Appendix.
3.3 Deterministic and probabilistic approaches. Loss functions.
As explained in the previous section, we have the option to use the ${\mathcal{M}}$ model to directly build the best prediction ${\widehat{{\bf{V}}}}_{t+h}$ or, using the output of ${\mathcal{M}}^{\prime}$, to compute it as the expectation under the best prediction of the probability law at each time according to Eq. (4). Let us describe these two options more precisely and the loss function one minimizes to estimate the parameters of ${\mathcal{M}}$ or ${\mathcal{M}}^{\prime}$ respectively.
3.3.1 Deterministic approach
Since it is generally not possible to compute the conditional mean square involved in Eq. (1) at each prediction step, the most common approach is to compute the vector of model parameters ${\bf{\Theta}}$ that minimizes the time-averaged value of ${\cal E}^{mse}_{d,t}$, the conditional MSE defined in Eq. (1). Empirically, this involves finding the parameters ${\bf{\Theta}}$ that minimize a loss function that corresponds to the empirical MSE defined as:
$\mathcal{E}^{mse}$ $=$ $N^{-1}\sum_{k=1}^{N}\left\|{\bf{V}}^{s}_{d,k}-{\widehat{{\bf{V}}}}^{s}_{d,k}\right\|^{2}$ (7)
$=$ $N^{-1}\sum_{k=1}^{N}\sum_{i=1}^{H}\left({\widehat{V}}^{s_{k}}_{d_{k},t_{k}+h_{i}}-V^{s_{k}}_{d_{k},t_{k}+h_{i}}\right)^{2}$ (8)
where $N$ is the number of available samples (in the training subset, see below) which corresponds to the number of different triplets “station”, “day” and “time of the day” ($s$,$d$,$t$) for which they are no missing data for building the observed velocity vector ${\bf{V}}^{s}_{d,t+h}$ and the input tensors defined in the previous section. In that case, the last layer illustrated in Fig. 2, is, as previously anticipated, a fully connected layer with the ReLu activation function and that directly outputs a 6-dimensional positive vector representing the values of ${\widehat{{\bf{V}}}}^{s}_{d,t+h}$ at 6 considered time horizon. The loss function we use is directly provided by Eq. (7).
3.3.2 Probabilistic approach
As mentioned previously, in addition to the previous deterministic forecast approach (also called “single-point”), which consists in finding the best parametrization of the conditional mean wind speed value, one can try to estimate directly the full conditional probability distribution from which the conditional mean is derived (see Eqs. (3), (4)).
Such an approach is usually referred to as probabilistic forecasting. It has become increasingly important in recent years as it captures not only the expected outcome but also uncertainties around this expectation and therefore provides notably insights into the likelihood of extreme events. A commonly used approach involves quantile-based methods which consists in directly estimating specific quantiles, such as the median or the 90th percentile, without assuming any particular probability distribution. On the other hand, parametric methods assume an underlying probability distribution and estimate its parameters to construct the forecast. These approaches are more parsimonious, offering a complete characterization of the forecast distribution, and they can integrate additional information, such as prior distributions, using frameworks like Bayesian inference. In this context, the challenge is to estimate, at each time step $t$, the vector of parameters ${\bf{\Pi}}_{t}$ characterizing the conditional law $f(y|{\bf{\Pi}}_{t})$. The most commonly used approach (see for example Salinas et al. (2020)) consists in considering the parameters that minimize the negative log-likelihood. Practically, this amounts to minimize the following loss function (also called logarithmic score):
$\ell=-N^{-1}\sum_{k=1}^{N}\ln f(V^{s_{k}}_{d_{k},t_{k}+h}|{\bf{\Pi}}_{t_{k}})$ (9)
where $V^{s_{k}}_{d_{k},t_{k}+h}$ is the observed surface wind speed at site $s_{k}$, day $d_{k}$ and time $t_{k}+h$. For a multi-horizon prediction, i.e. when the goal is to predict wind speed at horizons $h_{1},h_{2},\ldots,h_{H}$, $H\geq 1$ simultaneously, we can simply consider, as for the MSE loss in Eq. (7), the sum over all horizons of single horizon loss :
$\ell_{H}=-N^{-1}\sum_{k=1}^{N}\sum_{i=1}^{H}\ln f(V^{s_{k}}_{d_{k},t_{k}+h_{i}}|{\bf{\Pi}}_{t_{k}})\;.$ (10)
As discussed above, in this case the last layer of our neural network model ${\mathcal{M}}^{\prime}$ consists in a fully connected layer with output tensor of shape $(3,6)$, corresponding to the 3 M-Rice parameters at the six time horizons. For each set of parameters, the conditional mean ${\widehat{{\bf{V}}}}^{s}_{d,t+h}$ is then obtained as the mean value of the corresponding M-Rice law which is provided by Eq. (20) in Appendix.
@@ 3.3.2 Probabilistic approach
3.4 Training, validation and test sets, model’s hyper-parameters
As detailed in Sec. 2, the MeteoNet dataset covers a period of 3 years from 01-01-2016 to 31-12-2018. The input data in the ${\mathcal{M}}$, ${\mathcal{M}}^{\prime}$ models, is built from (normalized) data for each station $s$, day $d$ and time $t$ of the day ($t$ are expressed in multiple of hours and are such that $t\leq 18$ UTC). Let us refer to such a triplet ($s$, $d$, $t$) as a key. The stations we have considered in this work are the stations such that the number of available keys is greater than 2000, meaning that over all possible 17,520 keys for the whole period, at least 2000 are such that there is no missing data among all data presented in Table 1. Only 278 stations among the 478 available Meteonet stations satisfy this criterion and are the ones used in this study. This corresponds to a total number of keys, $N\simeq 3.5\times 10^{6}$. For learning purposes, this global sample is divided in 3 subsets: a training subset devoted to training with 70 % of data, a validation subset to monitor training and control the choice of parameters and hyper parameters of size 20 % of the total size and finally a test subset used to provide an unbiased assessment of the model’s performance with the remaining 10 % of the dataset. We remark that when on a given day $d$ and time $t$, data from a station $s$ are included in the training set, and data from the same station (or from a nearby station $s^{\prime}$) at a nearby time are included in the validation set, this can lead to overfitting. This occurs because similar situations appear in both the training and validation (or test) sets. To prevent this, the dataset is split by day: all data from a specific day are assigned exclusively to either the training, validation, or test set, ensuring no overlap across subsets.
@@ 3.4 Training, validation and test sets, model’s hyper-parameters
In this work, we will also experiment fine-tuning a model in order to optimize it for a specific station. Fine-tuning a neural network involves taking a pre-trained model and further training it on a specific subset of data, allowing it to adapt to localized conditions and improve predictive accuracy for a particular domain or task (Quinn et al., 2019) . Fine-tuning is often used when the general structure of the model is suitable but improvements are needed for specific data characteristics. It is different from the so-called “transfer learning” that refers to leveraging a model pre-trained on a large dataset and reusing its learned features for a new task. Transfer learning typically involves replacing or adding new layers to adapt the model to the new problem, while fine-tuning focuses on further optimizing the pre-trained model for better performance on the target dataset. To fine-tune the model for a specific station $s$, we initialize it with the parameters obtained from the global training phase. The fine-tuning process then involves further training the model using the same loss function, as defined in Eq. (10), but restricting the training data to that of station $s$ exclusively. To mitigate the risk of overfitting, the training data used for fine-tuning consist solely of the subset of the global training set corresponding to station $s$. Similarly, the validation and test sets are derived as subsets of the global validation and test datasets, containing only data associated with the station $s$.
Hyperparameters are parameters that govern the learning process and the configuration of the model. Unlike model parameters, they are not learned during the training phase, but are chosen before the learning process begins. While many hyperparameter optimization techniques that systematically explore configurations to improve model performance exist, such as grid search or random search, we did not apply such optimization methods in our work. Instead, we experimented with a few values for layer configurations and learning parameters.
The model configuration is shown in Fig. 2, and learning parameters are in Table 2. Since hyperparameters were not optimized on the validation set, we expect similar performance between validation and test sets. Our observations confirm this, with results within a few percentage points.
Table 2: Learning Hyperparameters used in our study
Parareter Value

### v2

@@ 4.1 Model inputs
The objective of this study is to explore the hybridization of NWP and observations using a streamlined neural network model. This “hybrid” neural network model that integrates NWP short-term forecasts with local ground station data is designed to process three distinct sets of MeteoNet input data: stations data, AROME predictions, and ARPEGE predictions. It is illustrated in Fig. 2 and precisely described in the next section. The model input feature variables ${\mathbf{X}}^{s}_{d,t}$ can be therefore defined as
${\mathbf{X}}^{s}_{d,t}=\Big[{\mathbf{GS}}^{s}_{d,t},{\mathbf{AR}}^{s}_{d,t},{\mathbf{AP}}^{s}_{d,t},{\mathbf{C}}^{s}_{d,t},{\mathbf{D}}^{s}\Big]$
where, as explained below, ${\mathbf{GS}}_{d,t}^{s}$${\mathbf{AR}}_{d,t}^{s}$ and ${\mathbf{AP}}_{d,t}^{s}$ are feature tensors built from respectively the ground station data, AROME and ARPEGE predictions at some given day $d$, time $t$ and corresponding to some given site $s$. The tensors ${\mathbf{C}}^{s}_{d,t}$ and ${\mathbf{D}}^{s}$ bring deterministic information like the day of the year, the time of the day, the location of the station and its neighboring stations (see below). Since NWP forecast data is only available for $0\leq t\leq 24$, the set of available features is constrained by the choice of the maximum forecast horizon, $h_{H}=6$ hours. Consequently, for each day $d$, the initial time $t$ is restricted to the range $t=0,1,\dots,17$ spanning the hourly intervals from 0:00 AM to 5:00 PM. It should also be noted that, prior to being input into the model, all tensors are standardized component-wise using their mean and variance computed from the training subset.
Each input component for a given location—that is, for any site referenced as a “ground station” in the MeteoNet dataset—is described more precisely below. These features are summarized in Table 1.
@@
Table 1: Summary of the MeteoNet input data used in the model $\mathcal{M}$} or $\mathcal{M}$$^{\prime}$
@@
First, ${\mathbf{GS}}_{d,t}^{s}$, associated with weather ground station data, is composed of 3 different variables: the wind speed data ($u$,$v$ components) and surface temperature data. Other meteorological variables, such as atmospheric pressure and humidity, exhibit a significantly lower availability of data. Consequently, we have chosen to focus exclusively on wind speed horizontal components and temperature as primary features. For each site, we incorporated these three variables as measured at the ten nearest surrounding weather stations where data were available, as well as at the site itself. Accordingly, at any given point defined by $(s,d,t)$, the total number of station-related input variables amounts to $N_{C}=33$. Furthermore, for each initial time $t$, we consider not only the observations at time $t$ but also those recorded at six preceding time steps, corresponding to t1t-1 through t6t-6, with each step representing a six-hour interval. It is important to note that the MeteoNet dataset provides ground station measurements at a temporal resolution of six minutes. However, our experiments have shown that increasing the time resolution does not lead to any significant improvement in forecasting performance within the considered model. Therefore, to maintain model simplicity and minimize the number of parameters, the analysis is restricted to data at an hourly scale. The number of time periods (past or present) is then $N_{H}=7$, so that the overall ${\mathbf{GS}}_{d,t}^{s}$ input tensor shape is $(7\times 33)$.
@@
In order to build ${\mathbf{AR}}_{d,t}^{s}$, among all AROME prediction data, we select wind eastward and northward components at 10 $m$, mean-sea-level pressure, 2 m temperature and relative humidity so that the number of input variables is $N_{C}=5$. We keep only data around the considered ground station position, i.e., we consider the set of $11\times 11$ AROME grid points centered around the ground station, namely if $i_{0},j_{0}$ are the coordinates of the grid point closest to the target ground station $s$, then the considered AROME predictions are those within a spatial extent of $\pm 0.125^{\circ}$ latitude/longitude around this grid point. Moreover, we keep all prediction horizons $t+1,t+2,\ldots,t+6$ so that $N_{H}=6$. The shape of the tensors ${\mathbf{AR}}^{s}_{d,t}$ is therefore $(6\times 5\times 11\times 11)$.
@@
The ${\mathbf{AP}}^{s}_{d,t}$ tensor, which contains the data from the ARPEGE NWP model, is constructed along the same line: we keep only wind eastward and wind northward components together with temperature at all 7 isobar levels and we also consider the pressure levels at all 7 heights. As far as prediction horizons are concerned, we consider the same 6 next hours set of horizons as for AROME. When the horizon is not available, we consider the closest horizon multiple of 3 hours (indeed, ARPEGE predictions are not available on a regular time grid, as the time step varies from 1 hour in the morning to 3 hours in the afternoon). Concerning the grid size, we use all grid points within an interval $\pm 0.2^{\circ}$ around the closest point to the station. Notice that this spatial extent is larger than former AROME one but, due to the lower resolution of ARPEGE, it corresponds to a smaller ($5\times 5$) grid size. Accordingly, the ARPEGE input tensors have a shape $(6\times 4\times 7\times 5\times 5)$.
Finally the tensors ${\mathbf{C}}^{s}_{d,t}$, contains a set of temporal and static features are also provided at station $s$, day $d$ and time $t$. As in, e.g., Baïle and Muzy (2023), the initial time $t$ and the day index $d$ are encoded using $\cos$ and $\sin$ functions, namely time $t$ is represented by $(\cos(\frac{2\pi t}{24}),\sin(\frac{2\pi t}{24}))$ and $d$ by $(\cos(\frac{2\pi d}{365}),\sin(\frac{2\pi d}{365}))$. Information about the location (latitude, longitude and altitude) of the station $s$ is also provided. An encoding of the position of each of the 10 neighboring sites of $s$, ${\mathbf{D}}^{s}$ with respect to $s$ is also used.
4.2 Model architecture
Figure 2: Schematic representation of the neural network architecture used to process station, AROME, and ARPEGE data. The output shape of each layer is indicated below the corresponding block. ”F.C.” denotes a fully connected layer, ”Flat.” stands for a flattening operation, ”Conc.” indicates a concatenation of features, and $LSTM^{*}$ refers to an LSTM layer that returns the full output sequence rather than only the final hidden state.
The overall model architecture, illustrated in Fig. 2, is composed of three branches devoted to handling the three different input features described previously. Rather than implementing a state-of-the-art architecture, this study adopts a simple neural network design that leverages conventional convolutional and recurrent layers for spatial and temporal feature extraction (Goodfellow et al., 2016). While the core architecture remains intentionally straightforward, we also experiment, as detailed in Section 4.4, with several architectural variants such as alternative temporal encoders and adjustments to layer configurations. Identifying more advanced architectures and optimizing model performance is left for future research.
@@ 4.2 Model architecture
One can see in Fig. 2 that the branches that process AROME data (${\mathbf{AR}}^{s}_{d,t}$) and ARPEGE data (${\mathbf{AP}}^{s}_{d,t}$) inputs begin with a time-distributed convolution (represented by blue 3D boxes), which applies the same spatial (2D for AROME and 3D for ARPEGE) convolution operation independently to each time step in a sequence of inputs, treating each time step as a separate sample. This allows the model to extract spatial features along the temporal axis while maintaining the temporal dimension for subsequent processing. For each of the $6$ times indices, the output of convolution layer is then flattened in order to provide a 2-dimensional tensor (red circular boxes on the sketch, of shape $(6,816)$ and ($6,208)$ for AROME and ARPEGE branches respectively), such that they have the same number of dimension as the Station branch (whose input shape is $(7,33)$). Each of the static tensors ${\mathbf{C}}^{s}_{d,t}$ and ${\mathbf{D}}^{s}$ are processed by a fully connected layer which encodes this information within a 16-dimensional tensor. These tensors are then repeated at each of the input times (7 times for station data and 6 times for AROME and ARPEGE branches) and then concatenated to former input tensors or outputs of convolution layers. The so-obtained tensors are then processed through 2 or 3 stacked “Long Short-Term Memory” (LSTM) layers, a recurrent neural network architecture that is widely used to process sequential information (Goodfellow et al., 2016) and which is represented by a green rectangle in Fig. 2. The outputs of all branches, as provided by a last fully connected layer, are then concatenated into a single tensor to be encoded by a final fully connected layer and then by the output layer.
As described previously, for the deterministic version of the model, $\mathcal{M}$} the output layer is, in agreement with Eq. (3) also a fully connected layer with a rectified linear unit (ReLU) activation that returns a vector of 6 positive values representing the components of ${\widehat{{\mathbf{V}}}}^{s}_{d,t+{\mathbf{H}}}$ the predicted wind speeds at horizons $h=1,2,\ldots,6$ hours. In the case of probabilistic prediction, as described by the model $\mathcal{M}$$^{\prime}$ in Eq. (6), the last layer is a fully connected layer which outputs a $(n,6)$ tensor, where $n=3$ is the number of parameters necessary to define the M-Rice probability distribution described in A.
@@
We will also consider the restriction of the model $\mathcal{M}$} to a single type of input, i.e., one that considers specifically $\Big[{\mathbf{GS}}^{s}_{t},{\mathbf{C}}^{s}_{d,t},{\mathbf{D}}^{s}\Big]$, $\Big[{\mathbf{AR}}^{s}_{t},{\mathbf{C}}^{s}_{d,t}\Big]$ or $\Big[{\mathbf{AP}}^{s}_{t},{\mathbf{C}}^{s}_{d,t}\Big]$ and their corresponding branch. When $\mathcal{M}$} is reduced to its “ground station”, “AROME” or its “ARPEGE” branch, the corresponding models will be denoted as respectively ${\mathcal{M}}_{GS}$ model, ${\mathcal{M}}_{AR}$ model and ${\mathcal{M}}_{AP}$ model.
Regarding the number of parameters to be learned in each model, $\mathcal{M}$$\mathcal{M}$$^{\prime}$, as previously described, consists of 262k parameters, while ${\mathcal{M}}_{AR}$, ${\mathcal{M}}_{AP}$, and ${\mathcal{M}}_{GS}$ contain 138k, 84k, and 39k parameters, respectively.
4.3 Baseline models
[…]
@@ 4.3 Baseline models
4.4 Training, validation and test sets, fine-tuning, model’s parameters and hyperparameters selection.
As detailed in Sec. 2, the MeteoNet dataset spans a three-year period from 1 January 2016 to 31 December 2018. The input data for the models $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$ are constructed from (normalized) values associated with each station $s$, day $d$, and initial time $t$ (with $t\leq 17:00$ UTC, expressed in hours). We refer to each triplet $(s,d,t)$ as a key. Stations included in this study are those with at least 2000 valid keys—i.e., at least 2000 triplets with no missing values among the features listed in Table 1. This filtering results in 278 stations (out of 478) being retained, yielding a total of approximately $N\simeq 3.5\times 10^{6}$ valid keys. To ensure reliable model training and evaluation, this global sample is divided into three mutually exclusive subsets: 70% for training, 20% for validation (used to monitor training and tune hyperparameters), and 10% for testing (used for final model evaluation). Importantly, to avoid potential temporal data leakage—i.e., the risk that highly correlated samples appear in both the training and validation/test sets—the partitioning is performed at the day level. That is, all keys from a given day $d$ are assigned exclusively to one of the three subsets, ensuring that no overlapping or closely correlated data points are shared across splits. This design prevents situations where, for example, data from station $s$ at day $d$ and time $t$ is in the training set, while data from the same station or a nearby one at the same or adjacent times appears in the validation or test set. Moreover, as detailed in Sec. 4.1, we only consider initial times up to 5:00 PM (17:00 UTC) each day. This additional constraint ensures that input sequences used for prediction on day $d$ do not incorporate data from the following day $d+1$ and remain sufficiently distant from the previous day d1d-1, further mitigating the risk of information leakage across subsets.
@@
In addition to global training, we also experiment with fine-tuning the model to optimize its performance for specific stations. Fine-tuning a neural network involves taking a pre-trained model and continuing its training on a given subset of data. This allows the model to adapt to a particular site, to capture local meteorological patterns and station-specific biases that may not be fully represented in the global training. This improves the predictive accuracy for a given domain or task (Quinn et al., 2019) which can be particularly beneficial in operational contexts. Fine-tuning is often used when the general structure of the model is suitable but improvements are needed for specific data characteristics. It is different from the so-called “transfer learning” that refers to leveraging a model pre-trained on a large dataset and reusing its learned features for a new task. Transfer learning typically involves replacing or adding new layers to adapt the model to the new problem, while fine-tuning focuses on further optimizing the pre-trained model for better performance on the target dataset. Unlike transfer learning—which typically involves modifying the model architecture to suit a new task—fine-tuning retains the original structure and focuses on refining the learned parameters for the target dataset. To fine-tune the model for a given station $s$, we initialize it with the parameters obtained from the global training phase. The model is then further trained using the same loss function defined in Eq. (7), but restricting the training data to samples associated with station $s$ only. To mitigate overfitting, the fine-tuning procedure maintains the same data partitioning as in the global training phase, but restricts the training, validation, and test sets to samples associated with the target station $s$.
Hyperparameters are settings that govern both the learning dynamics and architectural configuration of a neural network. Unlike model parameters, which are optimized during training, hyperparameters must be set prior to training. In this study, these include the spatial resolution of input data (e.g., AROME and ARPEGE grids), the number of units in LSTM and fully connected layers, the number of recurrent layers, dropout rates, batch sizes, and the learning rate of the optimizer. In this study, hyperparameter selection followed a two-stage strategy. First, we performed an initial manual exploration to establish a stable and reasonably well-performing baseline configuration. This initial phase combined the use of training hyperparameter values commonly adopted in the literature (e.g., typical values for learning rate, batch size, and dropout rate) with a manual exploration of various architectural alternatives to assess their impact on model performance. Specifically, we tested:
    Doubling the resolution of AROME and ARPEGE input data.
    Adjusting the number of units and layers for encoding static inputs (e.g., station position, time, and date).
    Halving or doubling the number of units in fully connected layers within each branch and in the final layers after merging.
    Modifying the number of LSTM layers or their hidden units.
    Substituting LSTM layers with alternative temporal encoders, such as GRU, Temporal Convolutional Networks (TCNs), or Transformer encoders, in the relevant branches.
None of these higher-level alternatives led to a consistent improvement in performance over the architecture presented in Section 4.1 (Fig. 2), which we retained for all reported results. In the second stage, we conducted a systematic grid search to fine-tune training hyperparameters around the baseline configuration. The grid search explored four key parameters: learning rate in $\{0.001,0.01\}$, batch size in $\{256,512,1024\}$, dropout rates (applied to all LSTM layers) in $\{0.01,0.02,0.2\}$ and the number of LSTM units in the largest LSTM layers $\{16,32,64\}$, with neighboring layers using half or a quarter of these values as indicated in Figure 2. Each model was trained for a fixed number of steps (specifically, $10$ epochs). The hyperparameter values selected through this process are reported in Table 2. It is noteworthy that the limited hyperparameter search space used to identify the best-performing configuration reduces the risk of overfitting to the validation set, thereby suggesting minimal differences between validation and test errors. This expectation is supported by the close agreement observed between validation and test performance, which differ by only a few percentage points—indicating robust generalization of the selected hyperparameters.
Table 2: Some of learning Hyperparameters used in this study
Parameter Value